# pFedGPA: Diffusion-based Generative Parameter Aggregation for Personalized Federated Learning

### v2

@@ Introduction
One challenge in FL is handling the heterogeneity of data distributions across various devices (Zhao et al. 2018; Kairouz et al. 2021). For example, in healthcare applications, patient data collected from different hospitals can vary greatly due to differences in patient demographics, medical equipment, and local practices. In linear aggregation methods such as FedAvg (McMahan et al. 2017a), data in different clients are assumed to share the same distribution. Non-IID data in local clients can lead to unstable results (Zhao et al. 2018; Karimireddy et al. 2020), as illustrated in Fig.  1. In this case, the parameters $\theta_{1}$ and $\theta_{2}$ from non-IID datasets are aggregated to form $\theta_{Fedavg}=(\theta_{1}+\theta_{2})/2$. While $\theta_{1}$ and $\theta_{2}$ are located in the high-probability regions of the parameter space and are well-optimized for their respective tasks, their average $\theta_{Fedavg}$ falls into a low-probability region, causing a collapse. This example highlights the limitations of linear aggregators and the need of more ingenious aggregation methods.
@@
To address the non-IID data challenge, several strategies for improving the global models have been proposed. FedProx (Li et al. 2020) introduced a local regularization term to optimize each client’s model.SCAFFOLD (Karimireddy et al., 2020b) introduces control variates to correct the drift in local updates. Clustering-based methods (Sattler, Müller, and Samek 2020; Li et al. 2020; Ghosh et al. 2020) cluster similar local models into groups and assign each group a global model. Moreover, self-supervised methods have been incorporated to define the similarity between model representations for correcting local training (Li, He, and Song 2021). Personalized FL (Smith et al. 2017) focuses on tailoring models for individual clients by combining information across clients. Decentralized MAML (Fallah, Mokhtari, and Ozdaglar 2020; Li et al. 2017) adapts the model-agnostic meta-learning framework to a federated setting, allowing clients to learn models through local adaptation (Finn, Abbeel, and Levine 2017; Nichol, Achiam, and Schulman 2018). Another category includes model mixing and layer adaptation strategies, where clients learn a mixture of global and local models (Deng, Kamani, and Mahdavi 2020; Hanzely and Richtárik 2020; Zhang et al. 2023). Decoupling model layers is also a popular and effective approach (Arivazhagan et al. 2019; Collins et al. 2021; Xu, Yan, and Huang 2022). For instance, FedRep (Collins et al. 2021) and FedBABU (Oh, Kim, and Yun 2022) train base layers globally using FedAvg, with personalized layers fine-tuned locally. Personalized aggregation based on model similarity has also been investigated (Zhang et al. 2021; Huang et al. 2021; Ye et al. 2023).
@@
Diffusion models have emerged as a powerful technique of generative AI, particularly excelling in producing high-quality images. Their superior properties have led to widespread applications in various vision (Lugmayr et al. 2022) and multi-modal tasks (Rombach et al. 2022; Kawar et al. 2023; Kumari et al. 2023).
The foundational work by (Sohl-Dickstein et al. 2015), studied non-equilibrium thermodynamics, highlighting its potential in generative modeling. (Ho, Jain, and Abbeel 2020) advanced this with Denoising Diffusion Probabilistic Models (DDPMs), improving the denoising process for better image synthesis. (Song and Ermon 2019) proposed Score-Based Generative Modeling, achieving remarkable results in diverse applications (Meng et al. 2021; Xu et al. 2022). (Song et al. 2020) further unified diffusion probabilistic modeling and score-based generative modeling using stochastic differential equations (SDEs), demonstrating their equivalence.
@@
Model parameter generation has progressed from basic gradient optimization to advanced meta-learning and Bayesian methods. Early efforts primarily focused on optimizing parameters using gradient-based techniques like Stochastic Gradient Descent (SGD) and its variants (Amari 1993). Meta-learning emerged to enable models to adapt quickly to new tasks with minimal data, exemplified by methods that use hypernetworks as metaknowledge to dynamically generate model parameters from input data (Zhmoginov, Sandler, and Vladymyrov 2022). Bayesian deep learnig used Variational Bayes to infer the distributions of model parameters, which are assumed to be Gaussian (Wilson and Izmailov 2020).
@@
For the heterogeneous data distribution setting, we adopt the approach proposed in (Karimireddy et al. 2020; Zhang et al. 2021), ensuring that all clients have equal data sizes. A portion of the data (s%, default 20%) is uniformly sampled from all classes, while the remaining (100 - s)is sampled from a set of dominant classes specific to each client. Clients are grouped based on their dominant classes, though this grouping is unknown to the server. Additionally, the size of the local training data is kept small, specifically at 600 samples per client, to emphasize the necessity of FL. The testing data on each client is drawn from the same distribution as the training data.

### v3

@@ Introduction
One challenge in FL is handling the heterogeneity of data distributions across various devices (Zhao et al. 2018; Kairouz et al. 2021). For example, in healthcare applications, patient data collected from different hospitals can vary greatly due to differences in patient demographics, medical equipment, and local practices. In linear aggregation methods such as FedAvg (McMahan et al. 2017a), data in different clients are assumed to share the same distribution. Non-IID data in local clients can lead to unstable results (Zhao et al. 2018; Karimireddy et al. 2020b), as illustrated in Fig.  1. In this case, the parameters $\theta_{1}$ and $\theta_{2}$ from non-IID datasets are aggregated to form $\theta_{Fedavg}=(\theta_{1}+\theta_{2})/2$. While $\theta_{1}$ and $\theta_{2}$ are located in the high-probability regions of the parameter space and are well-optimized for their respective tasks, their average $\theta_{Fedavg}$ falls into a low-probability region, causing a collapse. This example highlights the limitations of linear aggregators and the need of more ingenious aggregation methods.
@@
To address the non-IID data challenge, several strategies for improving the global models have been proposed. FedProx (Li et al. 2020) introduced a local regularization term to optimize each client’s model. SCAFFOLD (Karimireddy et al. 2020a) introduces control variates to correct the drift in local updates. Clustering-based methods (Sattler, Müller, and Samek 2020; Li et al. 2020) cluster similar local models into groups and assign each group a global model. Moreover, self-supervised methods have been incorporated to define the similarity between model representations for correcting local training (Li, He, and Song 2021). Personalized FL (Smith et al. 2017) focuses on tailoring models for individual clients by combining information across clients. Decentralized MAML (Fallah, Mokhtari, and Ozdaglar 2020) adapts the model-agnostic meta-learning framework to a federated setting, allowing clients to learn models through local adaptation (Finn, Abbeel, and Levine 2017; Nichol, Achiam, and Schulman 2018). Another category includes model mixing and layer adaptation strategies, where clients learn a mixture of global and local models (Hanzely and Richtárik 2020; Zhang et al. 2023). Decoupling model layers is also a popular and effective approach (Arivazhagan et al. 2019; Collins et al. 2021; Xu, Yan, and Huang 2022). For instance, FedRep (Collins et al. 2021) and FedBABU (Oh, Kim, and Yun 2022) train base layers globally using FedAvg, with personalized layers fine-tuned locally. Personalized aggregation based on model similarity has also been investigated (Zhang et al. 2021; Ye et al. 2023).
@@
Diffusion models have emerged as a powerful technique of generative AI, particularly excelling in producing high-quality images. Their superior properties have led to widespread applications in various vision (Lugmayr et al. 2022) and multi-modal tasks (Rombach et al. 2022; Kawar et al. 2023; Kumari et al. 2023). The foundational work by (Sohl-Dickstein et al. 2015), studied non-equilibrium thermodynamics, highlighting its potential in generative modeling. (Ho, Jain, and Abbeel 2020) advanced this with Denoising Diffusion Probabilistic Models (DDPMs), improving the denoising process for better image synthesis. (Song and Ermon 2019) proposed Score-Based Generative Modeling, achieving remarkable results in diverse applications (Meng et al. 2021; Xu et al. 2022). (Song et al. 2020) further unified diffusion probabilistic modeling and score-based generative modeling using stochastic differential equations (SDEs), demonstrating their equivalence.
@@
Model parameter generation has progressed from gradient optimization to advanced meta-learning and Bayesian methods. Early efforts primarily focused on optimizing parameters using gradient-based techniques like Stochastic Gradient Descent (SGD) and its variants (Amari 1993). Meta-learning emerged to enable models to adapt quickly to new tasks with minimal data, exemplified by methods that use hypernetworks as metaknowledge to dynamically generate model parameters from input data (Zhmoginov, Sandler, and Vladymyrov 2022). Bayesian deep learnig used Variational Bayes to infer the distributions of model parameters, which are assumed to be Gaussian (Wilson and Izmailov 2020).
@@
For the heterogeneous data distribution setting, we adopt the approach proposed in (Karimireddy et al. 2020b; Zhang et al. 2021), ensuring that all clients have equal data sizes. A portion of the data (s%, default 20%) is uniformly sampled from all classes, while the remaining (100 - s)is sampled from a set of dominant classes specific to each client. Clients are grouped based on their dominant classes, though this grouping is unknown to the server. Additionally, the size of the local training data is kept small, specifically at 600 samples per client, to emphasize the necessity of FL. The testing data on each client is drawn from the same distribution as the training data.